It has been projected by UNDP that the population of Wakanda will increase by 0.8%
for each of the next quarters for 5 years. If Wakanda registered a population of 25,000
in the second quarter of 2023 (base period):
a. What is the expected population in Wakanda after the third quarter in 2025?
b. What is the expected population in Wakanda after the first quarter in 2030?

Answer by Theo(13342)

growth rate of .8% per quarter gives a growth factor of 1.008 per quarter.
population in the second quarter of 2023 is 25,000.

a. What is the expected population in Wakanda after the third quarter in 2025?
population in second quarter of 2025 is 4 * (2025 minus 2023) = 4 * 2 = 8 quarters from second quarter of 2023.
population in third quarter of 2025 is 1 quarter more than that = 9 quarters from second quarter of 2023.
population in third quarter of 2025 will be 25,000 * 1.008 ^ 9 = 26858.68821.

b. What is the expected population in Wakanda after the first quarter in 2030?
second quarter in 2030 is 4 * (2030 - 2023) = 4 * 7 = 28 quarters from second quarter of 2023.
population in first quarter of 2030 is 1 quarter less than that = 27 quarters from second quarter of 2023.
population in first quarter of 2030 will be 25,000 * 1.008 ^ 27 = 31000.90524.

hee's what it looks like on a graph.

the equation graphed is y = 25000 * 1.008 ^ x